# A Simple and Effective Pruning Approach for Large Language Models

### v1

@@ 6 Related Work
Network Pruning and Sparsity. Pruning is a popular technique for compressing neural networks through the elimination of weights, yielding sparse networks (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Liu et al., 2019; Zhu et al., 2018). Pruning methods can be broadly categorized into structured and unstructured approaches. Structured pruning methods Xia et al. (2022b); Liu et al. (2017); Molchanov et al. (2017); Fan et al. (2020) remove entire structured components of a network, facilitating efficient GPU speedups, while unstructured methods like magnitude pruning operate at the individual weight level, maintaining performance even at higher sparsity levels. Existing pruning methods usually require either modifications to the training procedure Louizos et al. (2018); Sanh et al. (2020); Gale et al. (2019); Kusupati et al. (2020), retraining the pruned networks to regain accuracy Han et al. (2015, 2016); Liu et al. (2019), or an even more computationally intensive iterative retraining process Jonathan and Michael (2019); Zhu and Gupta (2017); Renda et al. (2020); Frankle et al. (2020). However, scaling these methods to LLMs with billions of parameters presents a challenge, as the required training process demands substantial computational resources Touvron et al. (2023); Zhang et al. (2022).
@@
Pruning with Limited Data. Most related to our approach is a recent line of work on pruning with limited data Hubara et al. (2021); Frantar et al. (2022); Frantar and Alistarh (2022); Kwon et al. (2022). Such methods require no modification to the original training procedure and also no retraining of the pruned networks on the full training dataset. The primary aim of these methods is to preserve performance during the pruning procedure, assuming access to a limited and small amount of data, also referred to as the calibration data. In order to mitigate the accuracy drop, a layer-wise reconstruction problem (Hubara et al., 2021) is solved to minimize the change of output evaluated on the calibration data. Existing popular solvers Frantar et al. (2022); Singh and Alistarh (2020) for the layer-wise reconstruction problem rely on heavy computation of second-order Hessian inverses, which do not scale to the large hidden state size of LLMs. SparseGPT (Frantar and Alistarh, 2023) develops an efficient weight update procedure for LLMs via synchronized second-order Hessian updates.
@@
Emergent Properties of LLMs. Our work is also related to recent studies on the existence of large magnitude outlier features in Transformer-based language models (Timkey and Schijndel, 2021; Bondarenko et al., 2021; Xiuying et al., 2022; Luo et al., 2021; Puccetti et al., 2022). Dettmers et al. (2022) demonstrates that when LLMs exceed a certain parameter scale (e.g. 6B), large magnitude features start to emerge and strongly affect all layers, which is considered as one of the emergent properties of LLMs (Dettmers et al., 2022; Wei et al., 2022a; Schaeffer et al., 2023). They further identify these emergent features as the cause of existing quantization methods’ failures. This observation has spurred the development of various quantization schemes Dettmers et al. (2022); Xiao et al. (2023); Lin et al. (2023); Dettmers et al. (2023) tailored specifically for LLMs to manage these outlier features. Our work extends this understanding, demonstrating that these emergent large magnitude features should also serve as pivotal indicators for determining which weights to prune in LLMs.
@@ 7 Conclusion
In this work, we propose a simple and effective method for pruning Large Language Models (LLMs). Inspired by the recent discovery of emergent large features in LLMs, our approach, termed Wanda (Pruning by Weights and activations), removes weights with the smallest magnitudes multiplied by the corresponding input activation norms, on a per-output basis. Without the need for any retraining or weight update procedures, Wanda is able to identify effective sparse networks within pretrained LLMs. Wanda could be used as a general pruning approach beyond LLMs. We hope our work contributes to a better understanding of sparsity in LLMs.
@@
Appendices
Appendix A Experimental Details
Our experiments are conducted on NVIDIA RTX A6000 GPUs with 50 GB memory. For pruning LLaMA models, we first load these models onto GPUs in 16-bit floating-point format. All subsequent procedures (e.g. pruning) are conducted directly on GPUs. For zero-shot tasks, we use the evaluation framework from https://github.com/EleutherAI/lm-evaluation-harness/.
Pruning Image Classifiers. We use the pretrained ConvNeXt-B model available at https://github.com/facebookresearch/ConvNeXt. For ViT, we use the DeiT-B model from https://github.com/facebookresearch/deit. The ConvNeXt-B and DeiT-B classifiers have 89M and 86M parameters respectively. As described in Section 5, we focus on pruning linear layers, which comprise the majority of the parameters in these two architectures. For ConvNeXt, the 11 convolutions in the its blocks can be viewed as linear layers, which account for approximately 95% of the total model parameters. For calibration data on ImageNet, we sample 4096 images from the training set, using them as the calibration set for our entire analysis. We find 4096 samples leads to a stable result for our pruning metric, beyond which we notice only marginal effect.
LoRA Fine-tuning. We adopt the LoRA fine-tuning code on LLaMA available at https://github.com/tloen/alpaca-lora. Our fine-tuning objective is the pretraining autoregressive language modeling objective. For the LoRA adapter, we update two weight matrices, and , in the self-attention module with a rank of 8. The fine-tuning process is conducted on the C4 training set, operating within a restricted computational budget that involves a single A6000 GPU and a time constraint of 5 hours.
Appendix B Additional Results
B.1 Pythia and OPT
While prior public LLM models may not be as powerful as LLaMA, we test the generality of our approach by evaluating it on two representative LLM families before the release of LLaMA, namely Pythia [Biderman et al., 2023] and OPT Zhang et al. [2022]. We select two models with approximately the same level of total parameters: Pythia-12b 111https://huggingface.co/EleutherAI/pythia-12b and OPT-13b 222https://huggingface.co/facebook/opt-13b. The results (unstructured sparsity) are shown in Table 7. We can see that magnitude pruning deteriorates significantly faster on Pythia-12b and OPT-13b in comparison to LLaMA models. For instance, even a sparsity ratio of 20% pushes the perplexity of pruned models to a meaningless level (greater than 1000). In contrast, our method Wanda consistently finds exact and effective sparse networks within both pretrained LLMs.
@@ B.1 Pythia and OPT
| Model | Dense | Pruning Method | Weight Update | 10% | 20% | 30% | 40% | 50% |

### v3

@@ 6 Related Work
Network Pruning and Sparsity. Pruning is a popular technique for compressing neural networks through the elimination of weights, yielding sparse networks (LeCun et al., 1989; Hassibi et al., 1993). It can be broadly categorized into structured and unstructured approaches.
Structured pruning methods (Liu et al., 2017; Molchanov et al., 2019; Fan et al., 2020; Shen et al., 2022; Xia et al., 2022; Fang et al., 2023; Nova et al., 2023), sometimes referred to as activation pruning (Gale et al., 2019; Dhillon et al., 2018), remove entire structured components of a network, facilitating efficient GPU speedups. Some existing methods (Babaeizadeh et al., 2016; Dubey et al., 2018) have explored structured pruning based on activation statistics of neuron/filter output, e.g. percentage of zero activations (Hu et al., 2016) and activation mean (Molchanov et al., 2017). Recently, Ma et al. (2023) have studied structured pruning of LLMs. Bansal et al. (2023); Liu et al. (2023b) and Elena Voita (2023) have demonstrated the existence of prompt-dependent and task-specific sparsity in the structural components of LLMs, e.g., attention heads and MLP neurons.
@@
Unstructured methods (Han et al., 2015; 2016; Paul et al., 2023; Hoang et al., 2023; Gadhikar et al., 2023; Liu et al., 2023a) like magnitude pruning operate at the individual weight level, maintaining performance even at higher sparsity levels. Existing pruning methods usually require either modifications to the training procedure (Sanh et al., 2020; Kusupati et al., 2020), retraining the pruned networks to regain accuracy (Liu et al., 2019; Zhou et al., 2023), or an even more computationally intensive iterative retraining process (Renda et al., 2020; Frankle et al., 2020). However, scaling these methods to LLMs with billions of parameters presents a challenge, as the required training process demands substantial computational resources (Hoffmann et al., 2022; Zhang et al., 2022).
@@
Pruning with Limited Data. Most related to our approach is a recent line of work on pruning with limited data (Hubara et al., 2021; Frantar et al., 2022; Frantar & Alistarh, 2022; Kwon et al., 2022). Such methods require no modification to the original training procedure and also no retraining of the pruned networks on the full training dataset. The primary aim of these methods is to preserve performance during the pruning procedure, assuming access to a limited and small amount of data, also referred to as the calibration data. In order to mitigate the accuracy drop, a layer-wise reconstruction problem (Hubara et al., 2021) is solved to minimize the change of output evaluated on the calibration data. Existing solvers (Singh & Alistarh, 2020; Frantar et al., 2022) for the layer-wise reconstruction problem rely on heavy computation of second-order Hessian inverses, which do not scale to the large hidden state size of LLMs. SparseGPT (Frantar & Alistarh, 2023) develops an efficient weight update procedure for LLMs via synchronized second-order Hessian updates.
@@
Emergent Properties of LLMs. Our work is also related to recent studies on the existence of large magnitude outlier features in large language models (Kovaleva et al., 2021; Bondarenko et al., 2021; Timkey & Schijndel, 2021; Luo et al., 2021; Puccetti et al., 2022; Wei et al., 2022b). Dettmers et al. (2022) demonstrate that when LLMs exceed a certain parameter scale (e.g., 6B), large magnitude features start to emerge and strongly affect all layers, which can be seen as an emergent property of LLMs (Dettmers et al., 2022; Wei et al., 2022a; Schaeffer et al., 2023). They also pinpoint these emerging features as the reason why existing quantization methods fail. This observation has spurred the development of various quantization schemes (Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2023; Dettmers et al., 2023; Behdin et al., 2023) tailored specifically for LLMs to handle outlier features. Our work extends this understanding, demonstrating that outlier features should also serve as pivotal indicators of which weights to prune in LLMs.
@@ 7 Conclusion
In this work, we propose a simple and effective method for pruning Large Language Models (LLMs). Inspired by the recent discovery of emergent large magnitude features in LLMs, our approach, termed Wanda (Pruning by Weights and activations), removes weights with the smallest magnitudes multiplied by the corresponding input activation norms, on a per-output basis. Without the need for any retraining or weight update procedures, Wanda is able to identify effective sparse networks within pretrained LLMs. We hope our work contributes to a better understanding of sparsity in LLMs. Last, considering the fast speed of pruning with Wanda, it would be interesting to investigate whether Wanda can be useful in the setting of sparse training (Evci et al., 2020; Peste et al., 2021; Kuznedelev et al., 2023; Benbaki et al., 2023; Frantar et al., 2023b), where pruning has to be conducted repeatedly and thus the pruning efficiency is critical.
@@
Appendix A Image classifiers
We study how Wanda would perform against magnitude pruning on tasks where the latter has been widely used. We conduct a study on ImageNet-1K (Deng et al., 2009), a standard image classification task where magnitude pruning has been extensively studied (Gale et al., 2019; Blalock et al., 2020). We consider two modern vision architectures: ConvNeXt (Liu et al., 2022) and Vision Transformer (ViT) (Dosovitskiy et al., 2021). We choose these two architectures mainly for two reasons: first, as LLMs are based on Transformers, we would like to test if our observations on LLMs still hold on Transformers for other tasks; second, as we are evaluating on image classification, we are interested in examining how these pruning methods work on ConvNet models, with ConvNeXt being a representative architecture.
We use two ImageNet-1K pretrained models: ConvNeXt-B and DeiT-B, with a top-1 accuracy of 83.8 and 81.8 respectively. We prune the linear layers only (for ConvNeXt, this includes equivalent 11 convolution layers). For calibration data, we sample 4096 images from ImageNet training set. We observe that 4096 samples lead to a stable result for our pruning metric, beyond which we notice only a marginal effect. We report the accuracy of one-shot pruned models without any subsequent retraining.
We first study whether pruning per output is superior over pruning per layer for pruning image classifiers. In Figure 3, we show comparison results for both the magnitude metric and the pruning metric of Wanda. We can see that for both ConvNeXt-B and DeiT-B, layer-wise pruning is slightly better than pruning per output. We then compare the pruning metric of Wanda and the magnitude metric on layer-wise pruning. Results are shown in Figure 4. Our novel pruning metric leads to better results than magnitude pruning, especially at high sparsities (e.g., 70 and 80).
Appendix B Wanda on previous LLMs
In addition to LLaMA and LLaMA-2, we experiment with three previous LLM model families: namely OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022) and Pythia (Biderman et al., 2023).
Comparison with Baselines. For OPT and Pythia, we experiment with varying sparsity levels (10 to 50). We conduct additional evaluation on OPT and BLOOM models with various sizes. Results are shown in Table 9, Table 10 and Table 11 respectively. Our observations are as follows:
Unlike LLaMA and LLaMA-2, the well-established magnitude pruning approach fails catastropically on OPT-13B and Pythia-12B, even for low sparsity levels (e.g., 20). This result further highlights the limitations of magnitude pruning for LLMs, as discussed in Section 3.
Unlike magnitude pruning, Wanda successfully prunes these LLMs to much higher sparsities across various LLM model families, without any weight update on the kept weights. This result shows that LLMs have effective sub-networks that are exact. We hope this observation could contribute to a better understanding of sparsity in LLMs.
There are cases where Wanda slightly underperforms SparseGPT, especially for OPT models (see Table 10), suggesting that for OPT, there may be a tradeoff between pruning speed and pruning accuracy. However, the gap between SparseGPT and Wanda tends to get smaller as model sizes increase. This can be seen in Table 10 and Table 11.
At lower sparsities (e.g., 20), Table 9 indicates that the computationally intensive weight update process may be unnecessary, as Wanda yields comparable or slightly superior results.
@@ Appendix B Wanda on previous LLMs
| Model | Dense | Pruning Method | Weight Update | 10 | 20 | 30 | 40 | 50 |
| OPT-13B | 10.13 | Magnitude | ✗ | 14.45 | 9e3 | 1e4 | 1e4 | 1e4 |
| SparseGPT | ✓ | 10.11 | 10.10 | 10.12 | 10.35 | 11.19 | ||
| Wanda | ✗ | 10.09 | 10.07 | 10.09 | 10.63 | 11.42 | ||